# references.bib

### v2

@@ I INTRODUCTION
The critical challenges of garment manipulation stem from the non-conformity of these objects with fundamental assumptions in robotics such as rigidity, known dynamics models, and low dimensional state space [longhini2024unfoldingliteraturereviewrobotic]. In the absence of such assumptions, state or partial-state estimation becomes an extremely high-dimensional prediction problem posed in partial observability conditions.
To circumvent this problem, some methods have abandoned state estimation and rely on task-specific extrinsic features. Wrinkles detection, for instance, has proven efficient for simple tasks such as cloth flattening [sun_heuristic_based], [seita2019deeptransferlearningpick]. These features can be homogeneously defined over the whole surface of the object as they are not dependent on specific object-attached points, which allows perception systems to locate them from purely geometrical information, such as point clouds or depth images, without occlusion reasoning. Furthermore, by default, they can be generalised to a wide range of objects of varying appearances. However, the absence of a notion of object state prevents their use in more general manipulation problems where understanding the object’s topology is crucial, e.g. dressing.
Others have relied on visual and visuo-spatial feature recognition by training deep neural networks to encode RGB or RGB-D observations in a constrained latent space [yan_learning_2020], [hoque_visuospatial_2021]. By using a full view of the object as input, these methods could address more complex manipulation tasks. However, learning these accurate and compressed latent representations for complex deformable objects requires a quantity of manipulation data that is presently unavailable.
Despite the technical limitations of inferring a relevant state representation from real-world observations, several works have proven, in simulation, the usefulness of extensive and reliable state information to achieve skills such as dressing or dynamic manipulation [clegg_learning_2018], [9981376], [10414107]. However, during training, these methods rely on information unavailable in real-world scenarios. Hence, they are either non-transferable or rely on perception transfer from simulated to real-world observations. The generation of perceptually convincing images for simulation is constrained by: 1) most simulated environments used for policy learning offer non-photorealistic rendering capabilities. 2) manually designing assets for high-fidelity rendering of real-life objects is a costly process requiring extensive human work.
@@ II-A Differentiable Rendering
With its impressive rendering accuracy, NeRF [mildenhall2021nerf] has been a cornerstone in establishing differentiable rendering as a relevant tool for multi-view scene capture. In parallel, DeepSDF [Park2019DeepSDFLC] has demonstrated the ability of deep neural architectures to accurately store an object’s geometry by fitting over point clouds. By combining these features, NeuS [wang2021neus] constrained NeRF’s underlying geometry representation to fit the definition of the signed distance function, thus enabling high-quality capture of both a scene’s geometry and appearance. While these methods mainly focused on rendering, they made it possible to integrate information from multiple points of view into a human-interpretable Euclidean 3D space. This architecture allows gradient flow to differentiable scene-embedded functions.
Object-NeRF [yang2021objectnerf], learnt distinct and composable object rendering functions for scene editing. PAC-NeRF [Li2023PACNeRFPA], PPR [yang2023ppr], and NDR [Cai2022NDR] paired differentiable rendering with physics-inspired deformation models producing coherent reconstruction over ordered sequences of images.
Differentiable rendering has also made its way into the robotics community, where it has been used for grasp detection in Grasp-NeRF [Dai2023GraspNeRF] and Dex-NeRF [IchnowskiAvigal2021DexNeRF]. But also as a 3D decoder in [Li20213DNS] for learning a predictive model for visuomotor control during fluid manipulation.
Finally, a similar approach has been used recently for data generation in the bin-picking context with COV-NeRF [mishra2024closing]. COV-NeRF introduces an object-composable NeRF that learns pickable objects’ geometrical and visual characteristics. Later, the objects are artificially re-arranged randomly to produce new and automatically labelled visuospatial data for downstream perception tasks. Our approach is similar in the general structure of the generation pipeline; however, because of the deformable nature of garments and the complex interaction they can have with their environment, we have relied on stronger shape priors by using triangle meshes, implying significant divergence between the methods.
@@ II-B Visual sim-to-real gap in cloth manipulation
While the gap between simulated data and real-world observation is a problem encountered in all simulation-based robotics, it poses an even greater challenge in textile manipulation as cloth simulators generally produce inaccurate results unless slow and compute-intensive methods, such as FEM, are used [blancomulero2024benchmarking]. The visual sim-to-real gap is a substantial problem resulting in significant performance degradation when transferring from simulation to real-world data, as exemplified by [8967827] or [wu_learning_2023]. The main problem underlying this gap is the challenge of state estimation in a crumpled configuration.
Some works have tried to address this problem using shape reconstruction [chi2021garmentnets], [MEDOR] from half-crumpled observations. They have demonstrated encouraging performance for state estimation.
In [SCRACT], the authors addressed the problem from the data generation angle by producing a dataset of observations with high-quality annotations. This is similar to the approach in this work, but we attain the same goal by automatic generation of convincing observations from known states rather than state estimation from observations with known action sequences.
Outside of robotic manipulation, realistic renderings of garments have been tackled in the virtual try-on community [bhatnagar2019mgn], [corona2021smplicit] with more recent works producing outstanding results [10147324]. Nonetheless, the challenge of generating realistic garment images in the context of try-ons is simplified by the human body constraining the garment shape in a low-occlusion configuration. To our knowledge, our work is the first to present a learnt automatic data generation pipeline for garments in any state.
@@ III-A COMPOSABLE SCENE RENDERING
Similar to [wang2021neus], we sub-divide the problem to represent the scene geometry on the one hand and, on the other hand, a colour function that attributes a colour to each point in the scene. More precisely, as shown in Fig. 2, we model the scene implicitly through three functions:
@@
[…]
With $w$ as the weighting function derived from the signed distance field of the scene from [wang2021neus].
@@ III-B MESH-BASED FIELDS
$\begin{split}\forall p\in\mathbb{R}^{3},p&=sdf_{\mathcal{M}}(p)\cdot n+q\\ \mathrm{With:}\quad q&=\operatorname*{arg\,min}_{u\in\mathcal{M}}(||p-u||_{2})\\ n&=\frac{p-q}{||p-q||_{2}}\end{split}$ (4)
@@
$\begin{split}\min_{X\in\mathbb{R}^{n\times k}}\sum_{ij}a_{ij}||X_{i}-X_{j}||_{2}^{2}&=\min_{X}tr(X^{t}LX)\\ s.t.\quad X^{t}DX&=I_{k}\end{split}$ (9)
With $A=(a_{ij})_{(i,j)\in\{1...n\}^{2}}$ the adjacency matrix of the graph, $D=diag(A)$ and $L=D-A$ the graph Laplacian.
As a result, this embedding preserves the notion of proximity between nodes. Additionally, it carries over meshes many of the intuitions of the more familiar sine-cosine embedding used in Euclidian spaces. For instance, it acts as a frequency separation operator on the mesh’s surface, allowing for faster learning of our feature fields. This behaviour can be observed in Fig. 3b, where the first coordinate displays two local optima, the second and third have three, and the fourth has four, thus increasing the amount of variation. Finally, LPE’s coordinates values are in the $[-1,1]$ interval, providing an adequately normalised input to downstream neural networks. During our training, we used an embedding size of $k=256$.
@@ III-C TRAINING
$\begin{split}\mathcal{L}_{depth}=\sum_{p\in\mathcal{D}_{s}^{c}}||\mathrm{S}(p)||_{2}\end{split}$ (11)
@@
$\begin{split}\mathcal{L}_{reg}=\sum_{x}(||\nabla\mathrm{S}(x)||_{2}-1)^{2}\end{split}$ (12)
@@
$\begin{split}\mathcal{L}_{mesh}=\sum_{x}||\mathrm{S}_{o}(x)-{sdf}_{\mathcal{M}}(x)||_{2}\end{split}$ (13)
@@ IV-A Rendering
Sock 0 $28.14\pm 2.52$ $0.82\pm 0.06$ $0.09\pm 0.04$
Sock 1 $25.48\pm 7.25$ $0.76\pm 0.18$ $0.07\pm 0.03$
Sock 2 $28.80\pm 3.06$ $0.83\pm 0.06$ $0.06\pm 0.02$
Sock 3 $27.48\pm 3.83$ $0.83\pm 0.08$ $0.06\pm 0.02$
@@ V CONCLUSIONS
In the future, we would like to further push the fidelity and ease of use of GARField by drawing inspiration, for instance, from the dynamic research community built around NeRF [mildenhall2021nerf].

### v3

# GARField: Addressing the visual Sim-to-Real gap in garment manipulation with mesh-attached radiance fields

@@ I INTRODUCTION
The critical challenges of garment manipulation stem from the non-conformity of these objects with fundamental assumptions in robotics such as rigidity, known dynamics models, and low dimensional state space [1]. In the absence of such assumptions, state or partial-state estimation becomes an extremely high-dimensional prediction problem posed in partial observability conditions.
To circumvent this problem, some methods have abandoned state estimation and rely on task-specific extrinsic features. Wrinkles detection, for instance, has proven efficient for simple tasks such as cloth flattening [2], [3]. These features can be homogeneously defined over the whole surface of the object as they are not dependent on specific object-attached points, which allows perception systems to locate them from purely geometrical information, such as point clouds or depth images, without occlusion reasoning. Furthermore, by default, they can be generalised to a wide range of objects of varying appearances. However, the absence of a notion of object state prevents their use in more general manipulation problems where understanding the object’s topology is crucial, e.g. dressing.
Others have relied on visual and visuo-spatial feature recognition by training deep neural networks to encode RGB or RGB-D observations in a constrained latent space [4], [5]. By using a full view of the object as input, these methods could address more complex manipulation tasks. However, learning these accurate and compressed latent representations for complex deformable objects requires a quantity of manipulation data that is presently unavailable.
Despite the technical limitations of inferring a relevant state representation from real-world observations, several works have proven, in simulation, the usefulness of extensive and reliable state information to achieve skills such as dressing or dynamic manipulation [6], [7], [8]. However, during training, these methods rely on information unavailable in real-world scenarios. Hence, they are either non-transferable or rely on perception transfer from simulated to real-world observations. The generation of perceptually convincing images for simulation is constrained by: 1) most simulated environments used for policy learning offer non-photorealistic rendering capabilities. 2) manually designing assets for high-fidelity rendering of real-life objects is a costly process requiring extensive human work.
@@ II-A Differentiable Rendering
With its impressive rendering accuracy, NeRF [9] has been a cornerstone in establishing differentiable rendering as a relevant tool for multi-view scene capture. In parallel, DeepSDF [10] has demonstrated the ability of deep neural architectures to accurately store an object’s geometry by fitting over point clouds. By combining these features, NeuS [11] constrained NeRF’s underlying geometry representation to fit the definition of the signed distance function, thus enabling high-quality capture of both a scene’s geometry and appearance. While these methods mainly focused on rendering, they made it possible to integrate information from multiple points of view into a human-interpretable Euclidean 3D space. This architecture allows gradient flow to differentiable scene-embedded functions.
Object-NeRF [12], learnt distinct and composable object rendering functions for scene editing. PAC-NeRF [13], PPR [14], and NDR [15] paired differentiable rendering with physics-inspired deformation models producing coherent reconstruction over ordered sequences of images.
Differentiable rendering has also made its way into the robotics community, where it has been used for grasp detection in Grasp-NeRF [16] and Dex-NeRF [17]. But also as a 3D decoder in [18] for learning a predictive model for visuomotor control during fluid manipulation.
Finally, a similar approach has been used recently for data generation in the bin-picking context with COV-NeRF [19]. COV-NeRF introduces an object-composable NeRF that learns pickable objects’ geometrical and visual characteristics. Later, the objects are artificially re-arranged randomly to produce new and automatically labelled visuospatial data for downstream perception tasks. Our approach is similar in the general structure of the generation pipeline; however, because of the deformable nature of garments and the complex interaction they can have with their environment, we have relied on stronger shape priors by using triangle meshes, implying significant divergence between the methods.
@@ II-B Visual sim-to-real gap in cloth manipulation
While the gap between simulated data and real-world observation is a problem encountered in all simulation-based robotics, it poses an even greater challenge in textile manipulation as cloth simulators generally produce inaccurate results unless slow and compute-intensive methods, such as FEM, are used [20]. The visual sim-to-real gap is a substantial problem resulting in significant performance degradation when transferring from simulation to real-world data, as exemplified by [21] or [22]. The main problem underlying this gap is the challenge of state estimation in a crumpled configuration.
Some works have tried to address this problem using shape reconstruction [23], [24] from half-crumpled observations. They have demonstrated encouraging performance for state estimation.
In [25], the authors addressed the problem from the data generation angle by producing a dataset of observations with high-quality annotations. This is similar to the approach in this work, but we attain the same goal by automatic generation of convincing observations from known states rather than state estimation from observations with known action sequences.
Outside of robotic manipulation, realistic renderings of garments have been tackled in the virtual try-on community [26], [27] with more recent works producing outstanding results [28]. Nonetheless, the challenge of generating realistic garment images in the context of try-ons is simplified by the human body constraining the garment shape in a low-occlusion configuration. To our knowledge, our work is the first to present a learnt automatic data generation pipeline for garments in any state.
@@ III-A COMPOSABLE SCENE RENDERING
Similar to [11], we sub-divide the problem to represent the scene geometry on the one hand and, on the other hand, a colour function that attributes a colour to each point in the scene. More precisely, as shown in Fig. 2, we model the scene implicitly through three functions:
@@
$\mathbf{d}(\mathbf{o},\mathbf{v})=\int_{0}^{+\infty}w_{S}(j)\cdot\mathrm{S}(\mathbf{p}(j))dj$ (3)
With $w$ as the weighting function derived from the signed distance field of the scene from [11].
@@ III-B MESH-BASED FIELDS
[…]
@@
[…]
With $A=(a_{ij})_{(i,j)\in\{1...n\}^{2}}$ the adjacency matrix of the graph, $D=diag(A)$ and L=DAL=D-A the graph Laplacian.
As a result, this embedding preserves the notion of proximity between nodes. Additionally, it carries over meshes many of the intuitions of the more familiar sine-cosine embedding used in Euclidian spaces. For instance, it acts as a frequency separation operator on the mesh’s surface, allowing for faster learning of our feature fields. This behaviour can be observed in Fig. 3b, where the first coordinate displays two local optima, the second and third have three, and the fourth has four, thus increasing the amount of variation. Finally, LPE’s coordinates values are in the [1,1][-1,1] interval, providing an adequately normalised input to downstream neural networks. During our training, we used an embedding size of $k=256$.
@@ III-C TRAINING
[…]
@@
[…]
@@
[…]
@@ IV-A Rendering
Sock 0 $28.14\pm 2.52$ $82\pm 0.06 0.0$$09\pm 0.04$
Sock 1 $25.48\pm 7.25$ $76\pm 0.18 0.0$$07\pm 0.03$
Sock 2 $28.80\pm 3.06$ $83\pm 0.06 0.0$$06\pm 0.02$
Sock 3 $27.48\pm 3.83$ $83\pm 0.08 0.0$$06\pm 0.02$
@@ V CONCLUSIONS
In the future, we would like to further push the fidelity and ease of use of GARField by drawing inspiration, for instance, from the dynamic research community built around NeRF [9].